OpenStudy (anonymous):

Sally has three pieces of material. The first piece is 1 yd. 2 ft. 6 in. long, the second piece is 2 yd. 1 ft. 5 in long, and the third piece is 4 yd. 2ft. 8in long. How much material does Sally have? A. 7 yd. 1 ft. 8 in. B. 8 yd. 4 ft. 4 in. C. 8 yd. 11 in. D. 9 yd. 7 in. E. 10 yd. Does anyone know how to calculate above?

OpenStudy (jlastino):

12 Inches = 1 foot 3 feet= 1 yard Add the number of inches first 6+5+8 = 19 From here we have 1 foot and 7 INCHES Add that 1 foot to the given 2+1+2+1 = 6 Converting it would give us 2 YARDS Add that to the given 2+1+2+4=9 So the answer is 9 yards and 7inches